(D³+4D²D¹+4DD¹^2)Z=0


1
Expert's answer
2021-06-22T08:37:55-0400

Since,highest power of Di.e., 3 is greater than highest power ofD'i.e., 2.

Therefore, put D=m and D'=1.

Then, the Auxiliary equation is,

m3+4m2+4m=0

m(m2+4m+4)=0

m(m+2)2=0

m=-2,-2,0

Therefore, the complementary function ia given by,

C.F= "\phi(y-2x)+x\phi(y-2x)+\phi(y)"